15 Tickets for a concert are given out randomly to a class with 20 students. No student is given more than 1 ticket.
i. find the number of ways in which this can be done.
There are 12 boys and 5 girls in the class. find the number of ways in which
i. 10 boys and 5 girls get the tickets.
ii. all the boys get the tickets.

Answer by stanbon(75887)

15 Tickets for a concert are given out randomly to a class with 20 students. No student is given more than 1 ticket.
i. find the number of ways in which this can be done.
Ans: 20C15 = 20C5 = (20*19*18*17*16)/(1*2*3*4*5) = 15504
--------------------
There are 12 boys and 5 girls in the class. find the number of ways in which
i. 10 boys and 5 girls get the tickets.
Ans: 12C10*5C5 = 12C2*5C5 = 66*1 = 66 ways
--------------------------------------------
ii. all the boys get the tickets.
Ans: 20C10 = 184756
----------------------
Cheers,
Stan H.
----------------------